# Demystifying Mergeability: Interpretable Properties to Predict Model Merging Success

### v1

@@ Abstract
Model merging combines knowledge from separately fine-tuned models, yet success factors remain poorly understood. While recent work treats mergeability as an intrinsic property, we show with an architecture-agnostic framework that it fundamentally depends on both the merging method and the partner tasks. Using linear optimization over a set of interpretable pairwise metrics (e.g., gradient $L_{2}$ distance), we uncover properties correlating with post-merge performance across four merging methods. We find substantial variation in success drivers (46.7% metric overlap; 55.3% sign agreement), revealing method-specific ”fingerprints”. Crucially, however, subspace overlap and gradient alignment metrics consistently emerge as foundational, method-agnostic prerequisites for compatibility. These findings provide a diagnostic foundation for understanding mergeability and motivate future fine-tuning strategies that explicitly encourage these properties.
@@ 6.3 Stable Metrics
While we just showed that optimal metric weightings are method-dependent, we now examine whether certain metrics exhibit stable predictive behavior across all merging methods. We define a stable metric as one whose average coefficient maintains a consistent sign across all four methods, indicating compatibility principles that transcend specific merging algorithms. To identify such metrics, we analyze the average coefficients $\bar{w}k$ and their standard deviations $\sigma_{k}$ over the $20$ LOTO folds for each method; metrics satisfying the sign-consistency criterion are reported in Table 4.
@@
Enhancing Mergeability during fine-tuning. To demonstrate the actionability of our findings, we conduct a toy experiment motivated by the importance of gradient distance. We fine-tune task-specific models with a gradient-magnitude penalty, encouraging smaller parameter updates and keeping models closer to the pretrained solution. Operating within a shared loss basin naturally promotes gradient alignment. Consistent with this intuition (Ortiz-Jimenez et al., 2024; Zhou et al., 2025a), the regularized models are more mergeable, yielding a modest but consistent improvement in post-merge performance ($\Delta=+0.55$). See Appendix A.5 for details. We emphasize that this experiment serves as a proof of concept; more principled fine-tuning strategies, targeting multiple signals and specific merging methods, are expected to yield larger gains.
@@ A.5.1 Motivation
Our coefficient analysis identified gradient-based metrics as important predictors of mergeability. In particular, the gradient distance between task-specific models correlates with post-merge performance degradation. This suggests that models whose gradients are more aligned, or equivalently, whose parameter updates from the pretrained initialization remain similar in direction and magnitude, may be more compatible for merging.
One simple way to encourage such alignment is to regularize the magnitude of parameter updates during finetuning. By penalizing large deviations from the pretrained weights, we bias all task-specific models to remain close to the shared initialization. This keeps the models within a similar region of the loss landscape, which naturally promotes gradient alignment and reduces gradient distance between any pair of finetuned models.
@@ A.5.2 Method
We augment the standard cross-entropy finetuning objective with an $L_{2}$ penalty on the parameter updates:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda\|\theta-\theta_{\text{pretrained}}\|_{2}^{2}$ (7)
where $\theta$ denotes the current model parameters, $\theta_{\text{pretrained}}$ is the pretrained initialization, and $\lambda$ controls the regularization strength.
This formulation is equivalent to $L_{2}$ regularization toward the pretrained weights rather than toward zero. It encourages the finetuned model to stay close to the pretrained solution while still adapting to the downstream task.

### v2

@@ Abstract
Model merging combines knowledge from separately fine-tuned models, yet success factors remain poorly understood. While recent work treats mergeability as an intrinsic property, we show with an architecture-agnostic framework that it fundamentally depends on both the merging method and the partner tasks. Using linear optimization over a set of interpretable pairwise metrics (e.g., gradient $L_{2}$ distance), we uncover properties correlating with post-merge performance across four merging methods. We find substantial variation in success drivers (46.7% metric overlap; 55.3% sign agreement), revealing method-specific “fingerprints”. Crucially, however, subspace overlap and gradient alignment metrics consistently emerge as foundational, method-agnostic prerequisites for compatibility. These findings provide a diagnostic foundation for understanding mergeability and motivate future fine-tuning strategies that explicitly encourage these properties.
@@ 6.3 Stable Metrics
While we just showed that optimal metric weightings are method-dependent, we now examine whether certain metrics exhibit stable predictive behavior across all merging methods. We define a stable metric as one whose average coefficient across the $20$ folds maintains a consistent sign across all four methods, indicating compatibility principles that transcend specific merging algorithms. To identify such metrics, we analyze the average coefficients $\bar{w}k$ over the $20$ LOTO folds for each method; metrics satisfying the sign-consistency criterion are reported in Table 4.
@@
Enhancing Mergeability during fine-tuning. To demonstrate the actionability of our findings, we conduct a toy experiment motivated by the importance of gradient distance. We fine-tune task-specific models with a gradient-magnitude penalty, encouraging smaller parameter steps and keeping models closer to the pretrained solution. Operating within a shared loss basin naturally promotes gradient alignment. Consistent with this intuition (Ortiz-Jimenez et al., 2024; Zhou et al., 2025a), the regularized models are more mergeable, yielding a modest but consistent improvement in post-merge performance ($\Delta=+0.55$). See Appendix A.5 for details. We emphasize that this experiment serves as a proof of concept; more principled fine-tuning strategies, targeting multiple signals and specific merging methods, are expected to yield larger gains.
@@ A.5.1 Motivation
Our coefficient analysis identified gradient-based metrics as important predictors of mergeability. In particular, the gradient distance between task-specific models correlates with post-merge performance degradation. This suggests that models whose gradients are more aligned may be more compatible for merging.
One way to encourage such alignment is to regularize gradient magnitudes during finetuning. By penalizing large gradients, we discourage sharp parameter updates that could push models into disparate regions of the loss landscape. This biases all task-specific models to update more gradually from the shared initialization, which naturally promotes gradient alignment and reduces gradient distance between any pair of finetuned models.
@@ A.5.2 Method
We augment the standard cross-entropy finetuning objective with a penalty on the squared magnitude of parameter gradients:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda\sum_{i}\|\nabla_{\theta_{i}}\mathcal{L}_{\text{CE}}\|_{2}^{2}$ (7)
where $\nabla_{\theta_{i}}\mathcal{L}_{\text{CE}}$ denotes the gradient of the cross-entropy loss with respect to parameter $\theta_{i}$, and $\lambda$ controls the regularization strength.
This formulation penalizes large gradient magnitudes during training, encouraging smaller and more moderate parameter updates at each step. By discouraging sharp updates, the finetuned models evolve more gradually from the pretrained initialization, remaining within similar regions of the loss landscape.